# Autodeleveraging: Impossibilities and Optimization

### v1

@@ D.1 Capped Pro-Rata Water-Filling
1:Winner equities $e=\{e_{1},\dots,e_{n}\}$, Effective caps $\beta=\{\beta_{1},\dots,\beta_{n}\}$, Target budget $B_{T}$
@@
19:  $R\leftarrow R-e_{p_{k}}$ $\triangleright$ User $p_{k}$ becomes fully capped
@@
D.2 Risk-Aware Pro-Rata Water-Filling
The RAP algorithm augments the capped procedure by prioritizing accounts according to their “cap-to-weight” ratios $\beta_{i}/w_{i}$.
@@ D.2 Risk-Aware Pro-Rata Water-Filling
1:Winner equities $e=\{e_{1},\dots,e_{n}\}$, Effective caps $\beta=\{\beta_{1},\dots,\beta_{n}\}$, Risk weights $w=\{w_{1},\dots,w_{n}\}$, Target budget $B_{T}$
@@
7:Compute ratios $r_{i}\leftarrow\beta_{i}/w_{i}$ for all $i$ (treat $w_{i}=0$ as $r_{i}=\infty$)
@@
20:  $W_{\mathrm{rem}}\leftarrow W_{\mathrm{rem}}-e_{p_{k}}w_{p_{k}}$ $\triangleright$ User $p_{k}$ becomes fully capped
@@
Appendix E Risk-Aware Pro-Rata (RAP)
E.1 Examples of Risk-Aware Pro-Rata (RAP) and Next Deficit
@@
E.2 RAP Optimality and Convex Dominance
@@
the capped reverse-waterfilling with weights $w_{t,i}=\rho(\lambda_{t,i})=\lambda_{t,i}\psi_{i}(1/\lambda_{t,i})$ minimizes $\delta_{t}(h)$ among all $h$ with $\sum_{i}e_{t,i}h_{t,i}=b_{t}$ and $0\leq h_{t,i}\leq H_{t,i}$.
@@
where $C_{h}$ is a constant independent of $h$ (can depend on $\lambda_{t,i}$). Maximizing $\sum_{i}\rho(\lambda_{t,i})\,e_{t,i}h_{t,i}$ under $\sum_{i}s_{t,i}=b_{t}$ and $0\leq s_{t,i}\leq e_{t,i}H_{t,i}$ is a fractional knapsack problem solved by sorting the values $\rho(\lambda_{t,i})$. The optimizer for this problem is reverse-waterfilling [BoydVandenberghe2004]:
$h_{t,i}=\min\{H_{t,i},\ \tau^{\star}_{t}\,w_{t,i}\},\qquad w_{t,i}=\rho(\lambda_{t,i}),$
with $\tau^{\star}_{t}$ set by $\sum_{i}e_{t,i}h_{t,i}=b_{t}$. This choice of $w$ minimizes $\delta_{t}$ among all weighted reverse-waterfilling rules. ∎
@@
We first note that theoretical results from the measure theoretic literature imply than RAP should provide Schur-convex dominance. RAP weights $w_{i}=\lambda_{T,i}g(\lambda_{T,i})$ can be interpreted as allocating budget proportional to a coherent, law-invariant risk measure. Specifically, let $\rho(\lambda)=\lambda\,\psi(1/\lambda)$ be a risk density with $\psi$ convex and nonincreasing. This corresponds to a spectral risk measure $\rho(X)=\int_{0}^{1}\mathrm{ES}_{\alpha}(X)\,d\mu(\alpha)$ via the Kusuoka representation [Kusuoka2001, Acerbi2002]. Choosing $g(\lambda)=\rho(\lambda)/\lambda$ aligns the RAP allocation with this spectral density. Known results from [Kusuoka2001] then imply RAP with a “more convex” risk density (in the Schur sense) will weakly Schur-dominate any other weighted pro-rata rule with a less concentrated density. Instead of utilizing such strong measure theoretic tools, we instead prove this directly below in an elementary manner.
Theorem E.2 (Constructive Schur–convex dominance).
@@
We split this proof into three steps. The first step is analogous to the proof of Proposition C.1, where we analyze the change to the weights on a piecewise constant set of intervals. The second step is to analyze how the residuals change with budget using the piecewise-constant representation. Given the change in residual, the final step uses the majorization inequality to show that the residual vector of RAP weakly submajorizes that of any other weighted pro-rata rule.
Step 1: Active set and parameterization. We first define the active set $A(b)$ as in (36). Any weighted reverse-waterfilling with weights $w$ admits the water-level form
$h^{(w)}_{T,i}(b)\;=\;\min\{\beta_{i},\ \tau^{(w)}(b)\,w_{i}\},$
where $\tau^{(w)}(b)$ is chosen so that $\sum_{i}e_{T,i}\,h^{(w)}_{T,i}(b)=b$. On an interval $b_{0},b_{1}]$ with fixed $A(b)$ we have
$\frac{db}{d\tau^{(w)}}=\sum_{j\in A(b)}e_{T,j}w_{j},\qquad\frac{dh^{(w)}_{T,i}}{db}\;=\;\begin{cases}\displaystyle\frac{w_{i}}{\sum_{j\in A(b)}e_{T,j}w_{j}},&i\in A(b),\\[4.30554pt] 0,&i\notin A(b).\end{cases}$
@@
$z_{T,i}(b)\;=\;\rho(\lambda_{T,i})\,e_{T,i}\,\bigl(1-h_{T,i}(b)\bigr).$
Then on $b_{0},b_{1}]$,
$\frac{dz^{(w)}_{T,i}}{db}\;=\;\begin{cases}\displaystyle-\frac{\rho(\lambda_{T,i})\,e_{T,i}\,w_{i}}{\sum_{j\in A(b)}e_{T,j}w_{j}},&i\in A(b),\\[4.30554pt] 0,&i\notin A(b).\end{cases}$
For $\mathrm{RAP}(g^{\star})$ we take $w_{i}=\rho(\lambda_{T,i})$, giving
$\frac{dz^{(\mathrm{RAP})}_{T,i}}{db}\;=\;\begin{cases}\displaystyle-\frac{\rho(\lambda_{T,i})^{2}\,e_{T,i}}{\sum_{j\in A(b)}e_{T,j}\rho(\lambda_{T,j})},&i\in A(b),\\[4.30554pt] 0,&i\notin A(b).\end{cases}$
Step 3: Majorization at each budget. On a fixed $A(b)$, sort indices by decreasing $\rho(\lambda_{T,i})$. By the hypothesis that the haircut share vector of $h^{(w)}$ on $A(b)$,
$\sigma^{(w)}_{i}(b)\ :=\ \frac{e_{T,i}w_{i}}{\sum_{j\in A(b)}e_{T,j}w_{j}},$
is no more concentrated on high-$\rho$ indices than the RAP share $\sigma^{(\mathrm{RAP})}_{i}(b)=\frac{e_{T,i}\rho(\lambda_{T,i})}{\sum_{j\in A(b)}e_{T,j}\rho(\lambda_{T,j})}$, the rearrangement/majorization inequality implies that for every $k$,
$\sum_{i\leq k}\!\frac{dz^{(\mathrm{RAP})}_{T,(i)}}{db}\;\leq\;\sum_{i\leq k}\!\frac{dz^{(w)}_{T,(i)}}{db}$
where $(i)$ denotes the order by decreasing $\rho$. Hence the instantaneous decrease vector under RAP weakly submajorizes that under $h^{(w)}$ on $b_{0},b_{1}]$. Integrating over $b$ preserves $\prec_{w}$ on the interval, and concatenating the finitely many intervals where $A(b)$ changes preserves the order overall: $z_{T}(\mathrm{RAP}(g^{\star}))\preceq_{w}z_{T}(h^{(w)})$. Schur–convexity then yields the separable convex loss comparison. ∎
E.3 Example of the Solvency vs. Long-Term Revenue Trade-Off
@@
Appendix F Stackelberg Control
@@
F.1 Opposing Schur orderings for time to solvency and LTV
We formalize the fundamental trade-off between aggressive debt reduction (safety) and trader fee retention (value), formalizing the example in Appendix E.3. Let $n_{t}=|\mathcal{W}_{t}|$. Given any feasible strategy $\pi$, write $z_{t}(\pi)\in{\mbox{\bf R}}_{+}^{n_{t}}$ for the vector of residual debts at time $t$ (sorted in decreasing order). Let $h_{t,i}^{\pi}\in[0,1]$ be the haircut fraction for agent $i$ with equity $e_{t,i}$, so that the haircut mass is $m_{t,i}(\pi)=h_{t,i}^{\pi}e_{t,i}$. We write $m_{t}(\pi)\in{\mbox{\bf R}}_{+}^{n_{t}}$ for the corresponding vector of haircut masses. Let $Z_{t}\in{\mbox{\bf R}}_{+}^{n_{t}}$ denote the equity shock at time $t$, following the notation of Section 7. These evolve componentwise as
@@ F.1 Opposing Schur orderings for time to solvency and LTV
Proposition F.1 (Solvency-Revenue Trade-off).
@@
We first give a one–step illustration of Proposition F.1 and then show how its hypotheses can fail for the literal queue versus capped pro–rata policies.
@@
Consider a single round with two winners and common initial residuals $z_{0}=(1,1)$. Let strategy $A$ use aggressive haircuts $M_{1}(A)=(1,1)$, fully clearing both accounts so that $z_{1}(A)=(0,0)$. Let strategy $B$ use milder haircuts $M_{1}(B)=(1,0)$, clearing only the first account and leaving $z_{1}(B)=(0,1)$. Then $z_{1}(A)\preceq_{w}z_{1}(B)$ and $M_{1}(A)\succeq_{w}M_{1}(B)$, so the hypotheses of Proposition F.1 hold with $T=1$. For a Schur–concave, coordinate-wise nonincreasing stage value such as
@@
Now consider the familiar one–round example with two winners, equities $e=(1,1)$, and budget $b=1$. Under capped pro–rata we have haircuts $h^{\mathrm{PR}}=(\tfrac{1}{2},\tfrac{1}{2})$, masses $m^{\mathrm{PR}}=(\tfrac{1}{2},\tfrac{1}{2})$, and residuals $z^{\mathrm{PR}}=(\tfrac{1}{2},\tfrac{1}{2})$. Under a queue policy we instead have $h^{\mathrm{Q}}=(1,0)$, so $m^{\mathrm{Q}}=(1,0)$ and $z^{\mathrm{Q}}=(0,1)$. In weak submajorization order, $z^{\mathrm{PR}}\prec_{w}z^{\mathrm{Q}}$ and likewise $m^{\mathrm{PR}}\prec_{w}m^{\mathrm{Q}}$. Thus no choice of labels $A,B$ can make pro–rata simultaneously satisfy the safety-dominance condition $z_{t}(A)\preceq_{w}z_{t}(B)$ and the “more cumulative haircuts” condition $M_{t}(A)\succeq_{w}M_{t}(B)$ in Proposition F.1. Moreover, for any Schur–concave, coordinate-wise nonincreasing $G$ (for instance $G(M)=-\|M\|_{2}^{2}$), we have $G(m^{\mathrm{Q}})\leq G(m^{\mathrm{PR}})$, so in this toy case the safer policy (pro–rata) also delivers higher user value. This shows that while Proposition F.1 captures a structural opposing–orders phenomenon, its hypotheses do not hold mechanically for every queue versus capped pro–rata comparison.
F.2 Regret Analysis
@@
When convex constraints $c_{t}(x)\leq 0$ are imposed, applying OMD to the Lagrangian guarantees the same $O(\sqrt{T})$ regret and $O(T^{-1/2})$ average constraint violation [Hazan2019].
We refer the interested reader to the standard reference [Hazan2019] for a detailed proof.
@@
Lemma F.3 (Convexity and subgradient).
@@
Lemma F.3 verifies the curvature and bounded-subgradient assumptions required by the master OMD regret theorem (Theorem F.2) hold, so we can directly plug $\tilde{L}_{t}$ into that framework.
@@
In the main text, the severity controller selects a scalar severity $\theta_{t}\in[0,\Theta_{t}]$ that determines the haircut budget via $b_{t}=\theta_{t}D_{t}$. Optimizing over $b_{t}$ or $\theta_{t}$ is equivalent; we keep $b_{t}$ as the decision variable in the appendix for notational convenience. The loss is $f_{t}(b)=\tilde{L}_{t}(b)$, which is convex by Lemma F.3. We use the Euclidean divergence $D(x\|y)=\tfrac{1}{2}(x-y)^{2}$, reducing OMD to projected gradient descent.
Corollary F.4 (Severity Regret).
@@
The domain diameter is $\bar{b}$. By Lemma F.3, subgradients correspond to dual variables bounded by the maximum weight $G$. The result follows directly from Theorem F.2. ∎
@@
Corollary F.5 (MDIC Regret).
@@
F.3 Stackelberg vs. Nash in a Two-Round ADL Game
@@
In simultaneous play, there are two pure Nash equilibria and we effectively have a Coordination Game [FudenbergTirole1991]:
@@
Proposition F.6 (Stackelberg Dominance).
@@
F.4 Follower Strategic Responses

### v3

@@ E.1 Capped Pro-Rata Water-Filling
1:Winner endowments $w=\{w_{1},\dots,w_{n}\}$ (under PNL-only, positive PNL), Effective caps $\beta=\{\beta_{1},\dots,\beta_{n}\}$, Target budget $B_{T}$
@@
19:  $R\leftarrow R-w_{p_{k}}$ $\triangleright$ User $p_{k}$ becomes fully capped
@@
E.2 Risk-Aware Pro-Rata Water-Filling
The RAP algorithm augments the capped procedure by prioritizing accounts according to their “cap-to-weight” ratios $\beta_{i}/\tilde{w}_{i}$, where $\tilde{w}_{i}$ are risk weights (not to be confused with the endowment $w_{i}$).
@@ E.2 Risk-Aware Pro-Rata Water-Filling
1:Winner endowments $w=\{w_{1},\dots,w_{n}\}$ (under PNL-only, positive PNL), Effective caps $\beta=\{\beta_{1},\dots,\beta_{n}\}$, Risk weights $\tilde{w}=\{\tilde{w}_{1},\dots,\tilde{w}_{n}\}$, Target budget $B_{T}$
@@
7:Compute ratios $r_{i}\leftarrow\beta_{i}/\tilde{w}_{i}$ for all $i$ (treat $\tilde{w}_{i}=0$ as $r_{i}=\infty$)
@@
20:  $W_{\mathrm{rem}}\leftarrow W_{\mathrm{rem}}-w_{p_{k}}\tilde{w}_{p_{k}}$ $\triangleright$ User $p_{k}$ becomes fully capped
@@
Appendix F Risk-Aware Pro-Rata (RAP)
F.1 Examples of Risk-Aware Pro-Rata (RAP) and Next Deficit
@@
F.2 RAP Optimality and Convex Dominance
@@
the capped reverse-waterfilling with risk weights $\tilde{w}_{t,i}=\rho(\lambda_{t,i})=\lambda_{t,i}\psi_{i}(1/\lambda_{t,i})$ minimizes $\delta_{t}(h)$ among all $h$ with $\sum_{i}w_{t,i}h_{t,i}=b_{t}$ and $0\leq h_{t,i}\leq H_{t,i}$, where $w_{t,i}$ is the haircutable endowment (under PNL-only, $w_{t,i}=(\mathrm{PNL}_{t,i})_{+}$).
@@
where $C_{h}$ is a constant independent of $h$ (can depend on $\lambda_{t,i}$). Note that $e_{t,i}=c_{t,i}+\mathrm{PNL}_{t,i}$ and under PNL-only haircuts, $w_{t,i}=(\mathrm{PNL}_{t,i})_{+}$. The gradient $\partial\delta_{t}/\partial x_{t,i}$ where $x_{t,i}=h_{t,i}w_{t,i}$ is $\rho(\lambda_{t,i})e_{t,i}$ (expressed in equity terms because insolvency is an equity concept). Maximizing $\sum_{i}\rho(\lambda_{t,i})\,e_{t,i}(x_{t,i}/w_{t,i})$ under $\sum_{i}x_{t,i}=b_{t}$ and $0\leq x_{t,i}\leq w_{t,i}H_{t,i}$ is a fractional knapsack problem solved by sorting the values $\rho(\lambda_{t,i})e_{t,i}/w_{t,i}$. The optimizer for this problem is reverse-waterfilling [29]:
$h_{t,i}=\min\{H_{t,i},\ \tau^{\star}_{t}\,\tilde{w}_{t,i}\},\qquad\tilde{w}_{t,i}=\rho(\lambda_{t,i}),$
with $\tau^{\star}_{t}$ set by $\sum_{i}w_{t,i}h_{t,i}=b_{t}$. This choice of $\tilde{w}$ minimizes $\delta_{t}$ among all weighted reverse-waterfilling rules. ∎
@@
We first note that theoretical results from the measure theoretic literature imply than RAP should provide Schur-convex dominance. RAP weights $w_{i}=\lambda_{T,i}g(\lambda_{T,i})$ can be interpreted as allocating budget proportional to a coherent, law-invariant risk measure. Specifically, let $\rho(\lambda)=\lambda\,\psi(1/\lambda)$ be a risk density with $\psi$ convex and nonincreasing. This corresponds to a spectral risk measure $\rho(X)=\int_{0}^{1}\mathrm{ES}_{\alpha}(X)\,d\mu(\alpha)$ via the Kusuoka representation [78, 2]. Choosing $g(\lambda)=\rho(\lambda)/\lambda$ aligns the RAP allocation with this spectral density. Known results from [78] then imply RAP with a “more convex” risk density (in the Schur sense) will weakly Schur-dominate any other weighted pro-rata rule with a less concentrated density. Instead of utilizing such strong measure theoretic tools, we instead prove this directly below in an elementary manner.
Theorem F.2 (Constructive Schur–convex dominance).
@@
We split this proof into three steps. The first step is analogous to the proof of Proposition D.4, where we analyze the change to the weights on a piecewise constant set of intervals. The second step is to analyze how the residuals change with budget using the piecewise-constant representation. Given the change in residual, the final step uses the majorization inequality to show that the residual vector of RAP weakly submajorizes that of any other weighted pro-rata rule.
Step 1: Active set and parameterization. We first define the active set $A(b)$ as in (40). Any weighted reverse-waterfilling with risk weights $\tilde{w}$ admits the water-level form
$h^{(\tilde{w})}_{T,i}(b)\;=\;\min\{\beta_{i},\ \tau^{(\tilde{w})}(b)\,\tilde{w}_{i}\},$
where $\tau^{(\tilde{w})}(b)$ is chosen so that $\sum_{i}w_{T,i}\,h^{(\tilde{w})}_{T,i}(b)=b$ (budget constraint in endowment space). On an interval $b_{0},b_{1}]$ with fixed $A(b)$ we have
$\frac{db}{d\tau^{(\tilde{w})}}=\sum_{j\in A(b)}w_{T,j}\tilde{w}_{j},\qquad\frac{dh^{(\tilde{w})}_{T,i}}{db}\;=\;\begin{cases}\displaystyle\frac{\tilde{w}_{i}}{\sum_{j\in A(b)}w_{T,j}\tilde{w}_{j}},&i\in A(b),\\[4.30554pt] 0,&i\notin A(b).\end{cases}$
@@
$z_{T,i}(b)\;=\;\rho(\lambda_{T,i})\,e_{T,i}\,\bigl(1-h_{T,i}(b)\bigr),$
where the equity $e_{T,i}$ appears because insolvency is an equity concept, but the control variable is the endowment $w_{T,i}$ via $h_{T,i}$. Then on $b_{0},b_{1}]$,
$\frac{dz^{(\tilde{w})}_{T,i}}{db}\;=\;\begin{cases}\displaystyle-\frac{\rho(\lambda_{T,i})\,e_{T,i}\,\tilde{w}_{i}}{\sum_{j\in A(b)}w_{T,j}\tilde{w}_{j}},&i\in A(b),\\[4.30554pt] 0,&i\notin A(b).\end{cases}$
For $\mathrm{RAP}(g^{\star})$ we take $\tilde{w}_{i}=\rho(\lambda_{T,i})$, giving
$\frac{dz^{(\mathrm{RAP})}_{T,i}}{db}\;=\;\begin{cases}\displaystyle-\frac{\rho(\lambda_{T,i})^{2}\,e_{T,i}}{\sum_{j\in A(b)}w_{T,j}\rho(\lambda_{T,j})},&i\in A(b),\\[4.30554pt] 0,&i\notin A(b).\end{cases}$
Step 3: Majorization at each budget. On a fixed $A(b)$, sort indices by decreasing $\rho(\lambda_{T,i})$. By the hypothesis that the haircut share vector of $h^{(\tilde{w})}$ on $A(b)$,
$\sigma^{(\tilde{w})}_{i}(b)\ :=\ \frac{w_{T,i}\tilde{w}_{i}}{\sum_{j\in A(b)}w_{T,j}\tilde{w}_{j}},$
is no more concentrated on high-$\rho$ indices than the RAP share $\sigma^{(\mathrm{RAP})}_{i}(b)=\frac{w_{T,i}\rho(\lambda_{T,i})}{\sum_{j\in A(b)}w_{T,j}\rho(\lambda_{T,j})}$, the rearrangement/majorization inequality implies that for every $k$,
$\sum_{i\leq k}\!\frac{dz^{(\mathrm{RAP})}_{T,(i)}}{db}\;\leq\;\sum_{i\leq k}\!\frac{dz^{(\tilde{w})}_{T,(i)}}{db}$
[…]
F.3 Example of the Solvency vs. Long-Term Revenue Trade-Off
@@
Appendix G Stackelberg Control
@@
G.1 Opposing Schur orderings for time to solvency and LTV
We formalize the fundamental trade-off between aggressive debt reduction (safety) and trader fee retention (value), formalizing the example in Appendix F.3. Let $n_{t}=|\mathcal{W}_{t}|$. Given any feasible strategy $\pi$, write $z_{t}(\pi)\in{\mbox{\bf R}}_{+}^{n_{t}}$ for the vector of residual debts at time $t$ (sorted in decreasing order). Let $h_{t,i}^{\pi}\in[0,1]$ be the haircut fraction for agent $i$ with equity $e_{t,i}$, so that the haircut mass is $m_{t,i}(\pi)=h_{t,i}^{\pi}e_{t,i}$. We write $m_{t}(\pi)\in{\mbox{\bf R}}_{+}^{n_{t}}$ for the corresponding vector of haircut masses. Let $Z_{t}\in{\mbox{\bf R}}_{+}^{n_{t}}$ denote the equity shock at time $t$, following the notation of Section 7. These evolve componentwise as
@@ G.1 Opposing Schur orderings for time to solvency and LTV
Proposition G.1 (Solvency-Revenue Trade-off).
@@
We first give a one–step illustration of Proposition G.1 and then show how its hypotheses can fail for the literal queue versus capped pro–rata policies.
@@
Consider a single round with two winners and common initial residuals $z_{0}=(1,1)$. Let strategy $A$ use aggressive haircuts $M_{1}(A)=(1,1)$, fully clearing both accounts so that $z_{1}(A)=(0,0)$. Let strategy $B$ use milder haircuts $M_{1}(B)=(1,0)$, clearing only the first account and leaving $z_{1}(B)=(0,1)$. Then $z_{1}(A)\preceq_{w}z_{1}(B)$ and $M_{1}(A)\succeq_{w}M_{1}(B)$, so the hypotheses of Proposition G.1 hold with $T=1$. For a Schur–concave, coordinate-wise nonincreasing stage value such as
@@
Now consider the familiar one–round example with two winners, equities $e=(1,1)$, and budget $b=1$. Under capped pro–rata we have haircuts $h^{\mathrm{PR}}=(\tfrac{1}{2},\tfrac{1}{2})$, masses $m^{\mathrm{PR}}=(\tfrac{1}{2},\tfrac{1}{2})$, and residuals $z^{\mathrm{PR}}=(\tfrac{1}{2},\tfrac{1}{2})$. Under a queue policy we instead have $h^{\mathrm{Q}}=(1,0)$, so $m^{\mathrm{Q}}=(1,0)$ and $z^{\mathrm{Q}}=(0,1)$. In weak submajorization order, $z^{\mathrm{PR}}\prec_{w}z^{\mathrm{Q}}$ and likewise $m^{\mathrm{PR}}\prec_{w}m^{\mathrm{Q}}$. Thus no choice of labels $A,B$ can make pro–rata simultaneously satisfy the safety-dominance condition $z_{t}(A)\preceq_{w}z_{t}(B)$ and the “more cumulative haircuts” condition $M_{t}(A)\succeq_{w}M_{t}(B)$ in Proposition G.1. Moreover, for any Schur–concave, coordinate-wise nonincreasing $G$ (for instance $G(M)=-\|M\|_{2}^{2}$), we have $G(m^{\mathrm{Q}})\leq G(m^{\mathrm{PR}})$, so in this toy case the safer policy (pro–rata) also delivers higher user value. This shows that while Proposition G.1 captures a structural opposing–orders phenomenon, its hypotheses do not hold mechanically for every queue versus capped pro–rata comparison.
G.2 Regret Analysis
@@
When convex constraints $c_{t}(x)\leq 0$ are imposed, applying OMD to the Lagrangian guarantees the same $O(\sqrt{T})$ regret and $O(T^{-1/2})$ average constraint violation [66].
We refer the interested reader to the standard reference [66] for a detailed proof.
@@
Lemma G.3 (Convexity and subgradient).
@@
Lemma G.3 verifies the curvature and bounded-subgradient assumptions required by the master OMD regret theorem (Theorem G.2) hold, so we can directly plug $\tilde{L}_{t}$ into that framework.
@@
In the main text, the severity controller selects a scalar severity $\theta_{t}\in[0,\Theta_{t}]$ that determines the haircut budget via $b_{t}=\theta_{t}D_{t}$. Optimizing over $b_{t}$ or $\theta_{t}$ is equivalent; we keep $b_{t}$ as the decision variable in the appendix for notational convenience. The loss is $f_{t}(b)=\tilde{L}_{t}(b)$, which is convex by Lemma G.3. We use the Euclidean divergence $D(x\|y)=\tfrac{1}{2}(x-y)^{2}$, reducing OMD to projected gradient descent.
Corollary G.4 (Severity Regret).
@@
The domain diameter is $\bar{b}$. By Lemma G.3, subgradients correspond to dual variables bounded by the maximum weight $G$. The result follows directly from Theorem G.2. ∎
@@
Corollary G.5 (MDIC Regret).
@@
Linear regret for static policies.
The following toy instance shows that any static severity $\bar{\theta}$ can incur $\Omega(T)$ regret. Consider two regimes $A$ (liquidity) and $B$ (stress) alternating every round, with loss
[…]
The per-round optimizer is $\theta^{\star}_{t}=0$ in $A$ and $\theta^{\star}_{t}=\bar{\theta}$ in $B$, so
$\sum_{t=1}^{T}f_{t}(\theta)-f_{t}(\theta^{\star}_{t})=\tfrac{T}{2}\theta+\tfrac{T}{2}(\bar{\theta}-\theta)=\tfrac{T}{2}\bar{\theta}=\Omega(T)$
for any fixed $\theta\in[0,\bar{\theta}]$. By Theorem G.2, OMD (and MDIC) with $\eta_{t}\propto 1/\sqrt{T}$ attains $O(\sqrt{T})$ regret on the same sequence.
G.3 Stackelberg vs. Nash in a Two-Round ADL Game
@@
In simultaneous play, there are two pure Nash equilibria and we effectively have a Coordination Game [57]:
@@
Proposition G.6 (Stackelberg Dominance).
@@
G.4 Follower Strategic Responses